# Decoupled Data Consistency with Diffusion Purification for Image Restoration

### v5

@@ Abstract
Diffusion models have recently gained traction as a powerful class of deep generative priors, excelling in a wide range of image restoration tasks due to their exceptional ability to model data distributions. To solve image restoration problems, many existing techniques achieve data consistency by incorporating additional likelihood gradient steps into the reverse sampling process of diffusion models. However, the additional gradient steps pose a challenge for real-world practical applications as they incur a large computational overhead, thereby increasing inference time. They also present additional difficulties when using accelerated diffusion model samplers, as the number of data consistency steps is limited by the number of reverse sampling steps. In this work, we propose a novel diffusion-based image restoration solver that addresses these issues by decoupling the reverse process from the data consistency steps. Our method involves alternating between a reconstruction phase to maintain data consistency and a refinement phase that enforces the prior via diffusion purification. Our approach demonstrates versatility, making it highly adaptable for efficient problem-solving in latent space. Additionally, it reduces the necessity for numerous sampling steps through the integration of consistency models. The efficacy of our approach is validated through comprehensive experiments across various image restoration tasks, including image denoising, deblurring, inpainting, and super-resolution.
Keywords: diffusion model, inverse problem, image restoration, consistency model, latent diffusion, diffusion purification
@@ 1 Introduction
Image restoration problems naturally arise in many computer vision and biomedical imaging applications, examples of which include image denoising, deblurring, inpainting, and super-resolution [1, 2, 3]. The objective of solving an image restoration problem, also known as an inverse problem, is to estimate an unknown signal $\bm{x}_{\star}\in\mathbb{R}^{n}$ from its corrupted or subsampled measurements $\bm{y}\in\mathbb{R}^{m}$ of the following form:
$\bm{y}\;=\;\mathcal{A}(\bm{x}_{\star})+\bm{n},$ (1)
where $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a (nonlinear) forward operator, and $\bm{n}\in\mathbb{R}^{m}$ is additive noise.
In practice, the number of observed measurements is often significantly smaller than the input dimension (i.e., $m\ll n$), making the problem highly ill-posed and resulting in infinitely many solutions that satisfy Equation (1). To deal with this challenge, a common approach is to regularize the solution via enforcing image priors. In the field of image restoration, traditional priors such as sparsity have been thoroughly explored [2]. However, recent advancements have shifted focus towards employing deep generative models as learned priors for image restoration. This shift largely came to fruition by the work of Bora et al. [4], who demonstrated that deep generative models could achieve superior image reconstruction quality with fewer number of measurements compared to traditional sparsity-based methods [4, 5, 6]. Following this groundbreaking study, there has been a surge of interest in investigating and exploring various deep generative models for solving image restoration tasks, including deep image priors [7], compressed sensing with generative models (CSGM) [4], and their subsequent developments [1].
More recently, there has been a lot of attention in using diffusion models as image priors, as diffusion models have shown to achieve superior image sampling quality compared to existing generative methods [8, 9, 10, 11]. However, unlike with other types of generative models, adapting diffusion models as image priors to solve inverse problems is not straightforward. This is because the sampling (or generative) process for diffusion models is generally formulated for sampling from an unconditional distribution $p(\bm{x})$, where $p(\bm{x})$ denotes the data distribution. In contrast, solving inverse problems involves sampling from the conditional distribution $p(\bm{x}|\bm{y})$ to ensure data consistency. To sample from this conditional distribution, we need access to the likelihood gradients $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$, which are intractable. Many earlier existing works tackle this issue with projection-type methods that project the unconditional sample $\bm{x}_{t}$ at each iteration $t$ onto the measurement subspace during the sampling process [9, 12, 13, 14, 15]. However, since the projection step is vulnerable to measurement noise and can only be applied to linear inverse problems, more recent works deal with this issue by directly approximating the conditional score function $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ throughout the sampling process under certain assumptions [16, 17, 18].
However, most (if not all) of these existing techniques encounter inherent challenges due to the coupled nature of data consistency and the sampling process. That is, the data consistency step usually takes place within the reverse sampling process of diffusion models, resulting in several fundamental limitations. Firstly, notice that the quality of the reconstructed image is highly dependent on the total number of data consistency steps that we perform. Since data consistency occurs within the reverse sampling process, the total number of data consistency steps can only be at most the total number of sampling steps. This severely limits the usage of accelerated diffusion model samplers, such as denoising diffusion implicit models (DDIMs) [11] and consistency models [19], which perform one-step sampling. Secondly, this coupling also makes it difficult to generalize existing methods to latent (or stable) diffusion, as recent works [20, 21] have demonstrated that simply adding the likelihood gradient for data consistency in the presence of latent diffusion models is insufficient for maintaining data consistency. This restriction hinders usability in large-scale image restoration tasks, where fast and accurate inference is essential.
@@
Figure 1: Illustration of the difference between our proposed method and existing posterior sampling techniques for image restoration. Our method decouples the data consistency step from the diffusion process, whereas existing methods incorporate data consistency into the reverse sampling process of diffusion models by taking an additional gradient step towards the measurement space.
Our Contributions.
In this study, we address these challenges by decoupling the data consistency and reverse sampling processes. Our algorithm involves a two-stage process that alternates between (i) generating initial estimations that are consistent with the measurements by minimizing the data fidelity objective and (ii) refining the initial estimations by incorporating the image prior via diffusion purification [22]. This effectively decouples the two processes, which consequently has two advantages over the state-of-the-art diffusion model methods for image restoration:
  • Framework Versatility. Compared to existing methods, our framework can easily incorporate any diffusion model in a plug-and-play fashion. We demonstrate this flexibility by presenting experimental results that effectively leverage both latent diffusion and consistency models for image restoration.
  • Improved Efficiency. We demonstrate that by decoupling these processes, we can significantly reduce inference time by more than 5×5$\times$5 ×, all while achieving state-of-the-art performance. This contribution is particularly notable because we can incorporate accelerated samplers without leading to insufficient data consistency.
We conduct thorough experiments to validate our method and its advantages across a wide range of image restoration problems, including super-resolution, inpainting, and image deblurring.
Throughout these tasks, we demonstrate that our method can outperform existing methods across several quantitative metrics and perform extensive ablation studies to highlight our approach.
Notation and Organization.
The remainder of this paper is organized as follows. In Section 2, we briefly discuss the essential preliminaries. Subsequently, we introduce our proposed framework in Section 3. Moreover, we further introduce acceleration techniques of our approach in Section 4. In Section 5, we validate the effectiveness of our algorithm with experiments on various image restoration problems. In Sections 6 and 7, we discuss the relationship of our work with prior art, limitation and future directions. All extra technical details are postponed to the appendices.
@@
We denote scalar (function) as regular lower-case letters (e.g. $s_{t},f(t)$), vectors (function) with bold lower-case letters (e.g. $\bm{x},\bm{s}(\bm{x}_{t},t)$). We use $[N]$ to denote the set $\{1,2,\ldots,N\}$, $\mathbb{P}(\cdot)$ to denote the probability, $\mathbb{E}\left[\cdot\right]$ to denote expectation, $||\cdot||_{2}$ to denote L2 norm, $\mathcal{N}(\cdot)$ to denote gaussian distribution. We use $\bm{x}_{t}$ to denote a sample $\bm{x}$ at time step $t$. We use $\bm{x}_{0}$ to denote a clean sample, and use $\bm{s}_{\bm{\theta}}$ and $\bm{s}_{\bm{\phi}}$ for the estimated score functions in the pixel and latent spaces, respectively (discussed in detail later).
2 Preliminaries
In this section, we first briefly review the basic fundamentals of diffusion models [23, 8, 9, 10, 24, 25] and then discuss how to adapt them for solving image restoration problems. Subsequently, we present the main ideas behind diffusion purification, which serve as the backbone of our proposed method.
@@ 2.1 Unconditional Diffusion Models
Diffusion models [23, 8, 9, 10] generate an image with two processes, generally referred to as forward and reverse processes. The forward process is to diffuse the data distribution $p_{0}(\bm{x})$ to a standard Gaussian $p_{T}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ by gradually injecting noise. In the variance-preserving (VP) setting [26], the forward process can be written as a stochastic differential equation (SDE) of the form
@@
where $\beta_{t}>0$ controls the speed of diffusion and $\bm{w}$ follows the standard Wiener process. To sample an image, we can employ the reverse SDE, which takes the form
@@
where $p_{t}(\bm{x}_{t})$ is the distribution of $\bm{x}_{t}$ at the time $t$, $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function, and $\bar{\bm{w}}$ follows the reverse Wiener process [27]. In practice, we estimate the score function using a deep neural network $\bm{s}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which is trained using the denoising score matching objective [28]:
$\underset{\bm{\theta}}{\min}\,\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}|\bm{x}_{0}}\left[\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{0t}(\bm{x}_{t}|\bm{x}_{0})\right\|_{2}^{2}\right],$
where $t\sim\mathcal{U}(0,1)$, $\bm{x}_{0}\sim p_{0}(\bm{x})$, and $\bm{x}_{t}|\bm{x}_{0}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0})$. In practice, both the forward and reverse SDE can be discretized, with
$\bm{x}_{t}$ $=\sqrt{1-\beta_{t}}\bm{x}_{t-1}+\sqrt{\beta_{t}}\bm{\epsilon}_{t-1},$
$\bm{x}_{t-1}$ $=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{t}+\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$.
2.2 Diffusion Models for Image Restoration Problems
@@
To solve image restoration problems with diffusion models, one of the most commonly used techniques is to modify the unconditional reverse sampling process in Equation 3 by replacing the unconditional score with a conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ based on $\bm{y}$. As such, the conditional reverse SDE has the following form
$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})\}\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}}.$
@@
Then, by Bayes rule, we can decompose the conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ into
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})+\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}).$
While the unconditional prior score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ can be obtained from a pre-trained diffusion model, the likelihood gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ is usually intractable.
To tackle this issue, recent works [16, 17, 18] resort to approximating the likelihood term with tractable surrogates, built upon the idea of diffusion posterior sampling (DPS) [16]. DPS approximates the likelihood gradient with
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})\approx\eta\nabla_{\bm{x}_{t}}\|\mathcal{A}(\hat{\bm{x}}_{0}(\bm{x}_{t}))-\bm{y}\|_{2}^{2},$ (4)
where $\eta>0$ can be viewed as a step size for controlling the data consistency strength and $\hat{\bm{x}}_{0}(\bm{x}_{t}):=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ is the posterior mean, which can be estimated through Tweedie’s formula [29, 30]:
$\hat{\bm{x}}_{0}(\bm{x}_{t})$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}))$
$\approx\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)),$
where $\bar{\alpha}_{T}=\prod_{t=1}^{T}(1-\beta_{t})$.
However, despite being one of the most recognized algorithms, DPS suffers from several limitations, as previously briefly illustrated. Firstly, the inference time of the reverse sampling process is significantly increased, as the likelihood estimation needs to backpropagate over the entire score function $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ for each iteration $t$ of the sampling process. Since the score function is generally parameterized as a large deep network, this process can be computationally challenging, especially for tasks involving high-dimensional images. Secondly, as the data consistency steps take place within the reverse sampling process, the number of gradient steps for data consistency is limited by the reverse sampling steps, which could lead to insufficient data consistency when using accelerated samplers such as consistency models [19]. Furthermore, the performance of DPS is sensitive to the step size $\eta$ due to the coupled data consistency and reverse sampling. When $\eta$ is small, the reconstructions will not be consistent with the measurements $\bm{y}$. On the other hand, a large $\eta$ may push $\bm{x}_{t}$ away from the noisy data manifold on which the diffusion model is trained, resulting in serious artifacts. Lastly, naively adapting DPS to latent diffusion models often fails due to the nonlinearity of the decoder. Since most diffusion-based image restoration solvers are built upon the idea of DPS, they all suffer from these fundamental issues, motivating the need to decouple the data consistency and reverse sampling processses.
Figure 2: An illustration of our overall framework. Our algorithm involves a total of $K$ iterations which alternate between (i) generating an initial reconstruction by solely optimizing the data fidelity objective and (ii) refining the image via diffusion purification.

### v6

# DCDP: Decoupled Data Consistency via Diffusion Purification for Solving General Inverse Problems

@@ Abstract
Diffusion models have emerged as powerful generative priors, excelling in solving inverse problems owing to their ability to model complex data distributions. Existing methods typically enforce data consistency by guiding the reverse sampling process with additional likelihood gradients. However, computing these gradients introduces significant computational overhead, prolongs inference time, and most importantly constrains compatibility with accelerated samplers. To overcome these limitations, we propose DCDP (Decoupled Data Consistency via Diffusion Purification), a two-stage diffusion-based framework for solving inverse problems that disentangles data consistency from reverse sampling steps. Specifically, DCDP iteratively alternates between the following two phases: (i) a reconstruction phase that enforces data consistency, followed by (ii) a refinement phase that applies diffusion purification to impose the learned generative prior. DCDP is broadly applicable to pixel space and latent space diffusion, and consistency models, achieving significant speedup in inference time compared to most state-of-the-art methods. We demonstrate its effectiveness and generality across a range of inverse problems, including image denoising, (nonlinear) deblurring, inpainting, and super-resolution.
Keywords.
Diffusion model, inverse problems, latent diffusion, diffusion purification, plug-and-play.
Originally released on arXiv in March 2024; revised in June 2025.
@@ 1 Introduction
Inverse problems naturally arise in many image processing and scientific applications, such as image denoising, deblurring, inpainting, super-resolution, and phase retrieval [1, 2, 3]. Mathematically, the goal of solving an inverse problem is to estimate an unknown signal $\bm{x}^{\star}\in\mathbb{R}^{n}$ from its subsampled (and possibly corrupted) measurements $\bm{y}\in\mathbb{R}^{m}$:
$\bm{y}\;=\;\mathcal{A}(\bm{x}^{\star})+\bm{n},$ (1)
where $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a linear or nonlinear forward operator and $\bm{n}\in\mathbb{R}^{m}$ is the additive noise. Typically, these inverse problems are highly ill-posed with $m\ll n$, resulting in infinitely many solutions that satisfy (1).
A common approach to address this challenge is to regularize the solution by enforcing certain data priors. In the past, many classical handcrafted priors such as sparsity or total variation have been investigated in the literature [2]. However, handcrafted priors are often overly simplistic, limiting their ability to capture the complex structures inherent in natural images. To deal with the challenge, deep generative models have emerged as better priors, significantly outperforming handcrafted priors in solving inverse problems. This paradigm shift was pioneered by recent seminal work such as [4], which showed that deep generative priors enable superior image reconstruction with fewer measurements than traditional sparsity-based methods; see also subsequent developments [1].
More recently, diffusion models have gained increasing attention as image priors [5, 6, 7, 8]. These models represent a novel class of deep generative models that are capable of producing high-quality samples thanks to their powerful ability to model complex data distributions. Specifically, diffusion models operate by learning to reverse a forward process that progressively adds Gaussian noise to the data till it becomes indistinguishable from pure noise. The key to performing the reverse process is learning the score functions $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$, i.e., the gradients of the log-density of the noise-mollified data $\bm{x}_{t}$, which can be approximated with deep networks trained with the denoising score matching objective [9].
When applied to solving inverse problems, the reverse process needs to be guided by additional likelihood gradients $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ to ensure consistency with the measurements $\bm{y}$. However, since the likelihood gradient $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ is generally intractable, existing methods either iteratively project $\bm{x}_{t}$ onto the measurement subspace [6, 10, 11, 12, 13], or approximate the gradient under simplified assumptions [14, 15, 16]. These approaches face fundamental challenges due to the coupling of data consistency and the reverse sampling process. Specifically, data consistency is typically enforced at each diffusion step, so that the total number of consistency updates is constrained by the number of reverse steps. As a result, achieving high reconstruction quality often requires a large number of diffusion steps, limiting compatibility with accelerated samplers such as Denoising Diffusion Implicit Models (DDIMs) [8] or consistency models (CMs) [17], which operate with fewer steps by improved designs. Moreover, this coupling complicates the extension to latent diffusion models: recent works [18, 19] demonstrate that naively incorporating the likelihood gradient fails to ensure data consistency in the latent space, posing a significant challenge for fast and accurate inference in high-dimensional inverse problems [20].
@@
Figure 1: Illustration of the difference between our proposed method and existing posterior sampling techniques for solving inverse problems. Our method decouples data consistency from the diffusion process, unlike existing approaches that enforce it via gradient steps during reverse sampling.
Our Contributions. In this study, we tackle the challenges of solving general inverse problems with diffusion models by introducing a new method termed Decoupled Data Consistency via Diffusion Purification (DCDP), which decouples data consistency steps from the reverse sampling process. As illustrated in Figure 1, our algorithm iteratively alternates between: (i) a reconstruction phase that enforces measurement consistency by minimizing a data fidelity objective, and (ii) a refinement phase that refines the estimate using a diffusion-based image prior via diffusion purification [21]. Importantly, the data consistency update is executed independently of the reverse diffusion process, offering two key advantages over existing diffusion-based inverse problem solvers:
  • Versatility. Our framework can seamlessly integrate a broad range of diffusion models in a plug-and-play fashion, including latent diffusion for improved memory efficiency and consistency models for accelerated sampling.
  • Efficiency. By decoupling these processes, we can significantly reduce the inference time (see Figure 7), while achieving comparable performance compared to state-of-the-art algorithms.
Empirically, we conducted comprehensive experiments validating our method on both linear (super-resolution, inpainting, and image deblurring) and nonlinear inverse problems (nonlinear deblurring) across various datasets including FFHQ [22], LSUN-Bedroom [23], and ImageNet [24], achieving competitive results over existing methods with a faster inference speed.
@@
Notations. Throughout this paper, we denote scalars (functions) with regular lower-case letters (e.g., $s_{t},f(t)$) and vectors (vector functions) with bold lowercase letters (e.g., $\bm{x},\bm{s}(\bm{x}_{t},t)$). We use $[N]$ to denote the set $\{1,2,\ldots,N\}$, $p(\cdot)$ to denote the probability, $\mathbb{E}\left[\cdot\right]$ to denote the expectation, $\|\cdot\|_{2}$ to denote the $\ell_{2}$ norm, and we use $\mathcal{N}(\cdot)$ to denote the Gaussian distribution. We use $\bm{x}_{t}$ to denote a sample $\bm{x}$ at time step $t$. We use $\bm{x}_{0}$ to denote a clean sample, and $\bm{s}_{\bm{\theta}}$ and $\bm{s}_{\bm{\phi}}$ for the estimated score functions in the pixel and latent spaces, respectively.
2 Background
@@ 2.1 Unconditional Diffusion Models
Diffusion models [25, 5, 6, 7] generate samples by reversing a pre-defined forward process. The forward process iteratively diffuses the data distribution $p_{0}(\bm{x})$ with Gaussian noise until a sufficiently large timestep $T$, so that the noise-mollified data distribution $p_{T}(\bm{x})$ becomes indistinguishable from an isotropic Gaussian distribution with covariance $\sigma^{2}_{T}\bm{I}$. In the variance-preserving (VP) setting [6], the forward process can be written as a stochastic differential equation (SDE) of the form
@@
where $\beta_{t}>0$ controls the speed of diffusion and $\bm{w}$ follows the standard Wiener process. To sample an image, we can start from $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\sigma^{2}_{T}\bm{I})$ and reverse the forward process (2). As shown in [6], the forward process (2) can be reversed with either a reverse SDE:
@@
where $p_{t}(\bm{x}_{t})$ is the intermediate noise-mollified distribution at time $t$, $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the score function, and $\bar{\bm{w}}$ is the standard Wiener process [26]. In practice, the score function is estimated using a deep neural network $\bm{s}_{\bm{\theta}}$ with parameter $\bm{\theta}$ trained using the weighted denoising score matching objective [9]:
$\underset{\bm{\theta}}{\min}\;\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}}\bigl{[}\lambda(t)\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{0t}(\bm{x}_{t}\mid\bm{x}_{0})\|_{2}^{2}\bigr{]},$
where $\lambda(t)$ is a positive weighting function, $t\sim\mathcal{U}(0,1)$, $\bm{x}_{0}\sim p_{0}(\bm{x})$, and $\bm{x}_{t}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0})$, where $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition kernel, i.e., the distribution of $\bm{x}_{t}$ induced by (2) given $\bm{x}_{0}$. In the VP setting, the transition kernel takes the following form:
$p_{0t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{x}_{0},(1-\alpha_{t})\bm{I}),$ (6)
[…]
2.2 Diffusion Posterior Sampling for Solving Inverse Problems
@@
To solve inverse problems using diffusion models, one of the most commonly used techniques is to modify the unconditional reverse sampling process in (3) by replacing the unconditional score with a conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ conditioned on measurement $\bm{y}$. As such, the conditional reverse SDE has the following form
$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}}.$ (7)
@@
Using Bayes rule, the conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ can be decomposed into the unconditional score and the likelihood gradient:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})}_{\text{unconditional score}}+\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})}_{\text{likelihood gradient}}.$ (8)
While the unconditional score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ can be approximated using a pre-trained diffusion model $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ (see Section 2.1), the likelihood gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ is computationally intractable in general. In the past, many existing methods focused on estimating the likelihood gradient under simplified assumptions [27, 14, 15, 16]. Among them, one of the most popular approaches is diffusion posterior sampling (DPS) [14], which approximates the likelihood gradient as
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})\approx\nabla_{\bm{x}_{t}}\|\mathcal{A}(\hat{\bm{x}}_{0}(\bm{x}_{t}))-\bm{y}\|_{2}^{2},$ (9)
where $\hat{\bm{x}}_{0}(\bm{x}_{t}):=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ is the posterior mean, which can be computed with the unconditional score function through Tweedie’s formula:
$\hat{\bm{x}}_{0}(\bm{x}_{t})$ $=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}))$ (10)
$\approx\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)).$ (11)
Note the $\ell_{2}$ norm in (9) is a direct consequence of assuming Gaussian measurement noise; if the noise follows a different distribution, the norm can be replaced accordingly. Although effective, the DPS-based method still suffers from several limitations: (i) the number of data consistency steps is limited by the number of reverse sampling steps, potentially resulting in insufficient measurement consistency when employing accelerated samplers, and (ii) computing the approximated likelihood gradient (9) requires backpropagating through the entire score function $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ at each timestep $t$, which is computationally expensive since $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ is typically parameterized by a large deep network. Moreover, adapting DPS to latent diffusion models (LDMs) [28] is highly nontrivial due to the nonlinear nature of their encoders and decoders [19, 18].